According to FRED, in the fourth quarter of 2014, the labor force was 2,394,000, employment equaled 2,254,500, and working-age population was 2,984,600 in New Zealand. Calculate the a. Labor force participation rate


1
Expert's answer
2021-01-07T10:53:41-0500

Labor force participation rate

"= \frac{Labor \; force}{Working-age \; population} \times 100 \\

= \frac{2394000}{2984600} \times 100 \\

= 80.21"

Answer: 80.21 %